There is a triangle formed by the intersection of the following lines:

r1 = [0,1,7]+s[15,6,-12]

r2 = [4,-4,8]+s[4,-5,1]

r3 = [4,-4,8]+s[4,-5,1]

Find the area of this triangle


1
Expert's answer
2022-03-15T14:29:12-0400

"r_1 = <0,1,7>+s_1<15,6,-12>\\

r_2 = <4,-4,8>+s_2<4,-5,1>\\

r_3 =<4,-4,8>+s_3<4,-5,1>\\
A=r_1\cap r_2\\
0+15s_1=4+4s_2\\
1+6s_1=-4-5s_2\\
7-12s_1=8+s_2\\
s_2=-12s_1-1\\
1+6s_1=-4+60s_1+5\\
s_1=0, s_2=-1\\
A(0,1,7)"

"r_2=r_3"

There is not a triangle formed by the intersection of the following lines